# Classification of $f(R)$ Theories Of Inflation

### v1

@@ 3.2 Scalars
We choose to work in the spatially-flat gauge, where the scalar perturbation of the metric are parametrized by
@@
Also in this case, $\sigma_{\text{s}}$ depends only on $v$ and $\beta_{\alpha}$. From the expressions of $\sigma_{\text{t,s}}$ we can derive their expansion in $\alpha$ and apply the results of [13].
@@ 4 Predictions
The arbitrary coefficients $C_{\text{t,s}}$ can be chosen such that $w^{\text{C}}_{\text{t}}(\eta(t))=u(t)$ and $w^{\text{C}}_{\text{s}}(\eta(t))=\mathcal{R}\equiv H\frac{\delta\phi}{\dot{\bar{R}}}$, where $\mathcal{R}$ is the curvature perturbation in the spatially-flat gauge. Indeed, the quantities $w^{\text{C}}_{\text{t}}$ and $w^{\text{C}}_{\text{s}}$ can only be proportional to $u$ and $\mathcal{R}$, respectively, being gauge-invariant and conserved on superhorizon scales. Comparing the redefinitions (3.24) and (3.35) with the expression (3.13) we find
@@
Then, using the equations (2.17) and (2.19), the tensor-to-scalar ratio and the spectral indices read
@@
It is interesting to note that, since $r$ and $n_{\text{t,s}}$ depend only on $v$, $\beta_{\alpha}$ and $Q_{0\text{t,s}}$, they are analytic [^1] From [13] we know that also $Q_{0\text{t,s}}$ are analytic in $b_{i}$. in $b_{i}$.
@@
These results are formally the same found in [13] for single-field inflation in the Einstein frame.
We can now compare the results with data. The most up-to-date experimental constraints are obtained by combining the results from Planck 2018 collaboration [18], the dark energy spectroscopic instrument [19] and recent data from the Atacama Cosmology Telescope [20, 21]. The result gives
@@
for $k_{*}=0.05\ \text{Mpc}^{-1}$. From now on every quantity with the subscript “*” is understood to be at $k=k_{*}$. From (4.13) we can derive $\alpha_{*}$ for each class, plug it back in the leading order of the tensor-to-scalar ratio $r_{*}$. In this way we obtain a relation between $r_{*}$ and $n_{\text{s}*}$. We find
@@
Using the central value for $n_{\text{s}*}$ and taking into account the bound for the tensor-to-scalar ratio $r_{*}<0.035$ from the Bicep/KECK collaboration [22], we find that the accepted values for the first nonvanishing coefficient of the beta functions are
@@
In the literature, the number of e-folds before the end of inflation $N_{k}$ is often used as primary constraint, since there are various theoretical estimates that place it between 50 and 60. However, those derivations depend on the inflationary model, as well as on the reheating temperature (see [23] for a clear derivation of the estimate and [24] for a recent study). Moreover, the number of e-folds is also a frame-dependent quantity. For these reasons, we prefer to fix the free parameters of the models by means of the experimental data for $n_{\text{s}}$ and derive the prediction for $N_{k}$, instead of imposing $N_{k}$ in a range and then check the prediction for $n_{\text{s}}$. Then the reader can make their own interpretations of the results using the values for $N_{k}$ that they consider acceptable.
@@
If we accept that $N_{*}$ can be slightly larger than 60, then any $f(R)$ in class I is in agreement with data, provided that $b_{0}\lesssim-0.48$. On the other hand, if we impose $50\leq N_{*}\leq 60$ then class I is almost out of the two-sigma range. However, this is a mild tension and it is not enough to exclude the models in class I, considering that there are also some tensions between the data themselves [25, 24]. Functions of class II are in good agreement at one-sigma range.
@@ 5 Derivation of $f(R)$ from $\beta_{\alpha}$
In this section we follow the strategy explained in section 2 to obtain the expansions of all the background quantities for class I and II. In this way we can identify the general expression for the $f(R)$ that belong to those classes. Finally, we use the results of [13] to work out the power spectra.
@@
Using the expression for $v$ we can write the expansions for (3.25) and (3.36), which are also analytic in $b_{i}$
@@
From the coefficients of $\sigma_{\text{t,s}}$ we can directly obtain the solutions of the Mukhanov-Sasaki equation in the two cases using [13]. The results up to the order $\mathcal{O}(\alpha_{k}^{3})$ for tensors and $\mathcal{O}(\alpha_{k}^{2})$ for scalars are
@@
[…]
@@ 5.1 Class I
where $F_{0}$ and $H_{0}$ are constants with $[F_{0}]=0$ and $[H_{0}]=1$. Using (2.21) we can write the expansion for $R$ and then invert it. We find
@@
We see that it is possible to obtain $f(R)$ with inverse and/or fractional powers of $\tilde{R}$. However, those models can only be phenomenological ones, since they do not have a well defined Minkowski limit and/or they could not be used on every spacetime because $\tilde{R}$ could be negative in general.
In order to have more usable theories, we impose some constraints on the function. First, there should be no singularity for $\tilde{R}=0$, in order to have a well-defined Minkowski limit, as well as any other singularity for real $\tilde{R}$. Then we also avoid terms involving $\sqrt{\tilde{R}}$. These considerations force $n$ to be a positive integer [^2] A priori we could consider $n\in\mathbb{Q}$ and then fix the relations between the $b_{i}$ such that fractional powers of $R$ do not appear. However, this is not possible since the term proportional to $\tilde{R}^{-n}$ in the square bracket of (5.12) cannot be set to zero.. Such constraints are useful to obtain functions that are suitable in every spacetime background. However, if we are interested in functions that are guaranteed to be well defined around de Sitter spacetime only, we can relax those assumptions (see at the end of this subsection for an example).
@@
The first is Starobinsky model, while the second is pure quadratic gravity. In both cases a cosmological constant is present. However, in the latter it cannot be set to zero. Moreover, the $R^{2}$ term has the wrong sign and the theory propagates a ghost. This can be fixed by choosing a negative $F_{0}$, which would then leads to a negative cosmological constant. Therefore, the second model is not physically acceptable and we are left only with the Starobinsky one.
@@
Finally, if we relax one of the constraints on $f(R)$ and allow for terms containing powers of $\sqrt{\tilde{R}}$ we can have modifications to Starobinsky model, such as
@@
where $[\lambda]=0$. This model was studied in [26].
Using formulas from [13], we obtain the following expressions for power spectra and spectral indices
@@
The coefficient $C_{\text{t,s}}$ can be obtained from (4.2) and using the explict expressions for $F$, $H$ and $J_{\text{t,s}}$. We find
@@
This result is the same for every class. Note that the relation between $C_{\text{s}}$ and $C_{\text{t}}$ is the same as in single-field inflation with any potential [13].
@@ 5.2 Class II
In general, the higher the class the more complicated are the singularities in the expressions of $F$ and $H$ at $\alpha=0$. Indeed, choosing $b_{0}=0$ and $b_{1}\neq 0$, the solutions of the equations (2.19) are
@@
which have essential singularities. In this case inverting $R(\alpha)$ gives an expansion in powers of $1/\mathcal{W}_{0}$, where $\mathcal{W}_{0}$ is the first branch of Lambert function, defined as
@@
$f(R)=-\frac{6F_{0}H_{0}^{2}\tilde{R}^{2}}{(3n\ln\tilde{R})^{1/n}}\left[1-\frac{2}{3n^{2}\ln\tilde{R}}+\frac{24-36n+39n^{2}-b_{3}}{108n^{4}\ln^{2}\tilde{R}}+\mathcal{O}(\ln^{-3}\tilde{R})\right].$ (5.28)
@@ 6 Change of frame
In order to match the definition of $\tilde{\alpha}$ with the one used in the Einstein frame in [13] we choose
@@
becomes regular in $\alpha=0$. Therefore, in class I also $\tilde{H}(\tilde{\alpha})$ is regular in $\tilde{\alpha}=0$. In fact, we have
@@
which matches with the one obtained in [13]. Moreover, the expression for $\varepsilon$ in the Einstein frame is exact and gives
@@ 7 Class 0: other polynomial $f(R)$
We conclude by discussing a special class of theories. The reader might have noticed that the simplest $f(R)$ we can think of, i.e. polynomials, do not appear in any of the classes studied above, with the exception for the Starobinsky model and pure quadratic gravity in class I. The reason is that simple polynomial extensions of Starobinsky or Hilbert actions have a linear beta function. It is not possible to export the techniques from asymptotically-free quantum field theory and apply our RG analogy at the present stage. In particular, there are no leading logarithms that can be resummed into the running slow-roll parameter. Moreover, as discussed in section 2, in the assumption $\alpha\ll 1$ a linear beta function is sufficient to guarantee $\varepsilon\ll 1$ but leads to a constant $\eta_{H}$ [plus corrections of order $\mathcal{O}(\alpha)$] and slow-roll inflation cannot be realized. Nevertheless, this class still admits a systematic expansion in powers of $\alpha$ for the background quantities, since we only require $\alpha\ll 1$ to perturbatively solve the equations. In other words, theories with a linear beta function can still produce a phase of accelerated expansion, but the slow-roll condition $|\eta_{H}|\ll 1$ is violated, at least far enough in the past (see the discussion below).
@@
With appropriate choices of the coefficients we can obtain any polynomial extensions of Starobinsky model, such as
@@
and so on. The same $f(R)$ without the $R^{2}$ term can also be obtained, as well as Starobinsky model with the “wrong sign” in front of the $R^{2}$ term plus a positive cosmological constant, i.e.
@@
However, such theory propagates a scalar ghost and is not physically acceptable.
The result just found shows an important fact about Starobinsky model. Among all the polynomial $f(R)$ extensions of general relativity, Starobinsky model is the only one that allows for slow-roll inflation during the initial phase of the evolution of the universe. Moreover, since all the other polynomial $f(R)$ belong to a different class, calculations cannot be obtained by suitably modifying those of Starobinsky model, not even perturbatively. This can be seen from the fact that, like in the other classes, $F$ and $H$ are not analytic in the first nonvanishing coefficient of the beta function, in this case $\bar{b}$.
Besides their simple form, polynomials are particularly important because, in general relativity, they are generated by renormalization (together with all other possible scalar quantities build with the Riemann tensor, its covariant derivatives and the metric tensor). This means that they are always present and even if we consider them as small corrections, they cannot be viewed as such, at least in the context of slow-roll inflation. The only way to avoid their radiative generation is to consider renormalizable theories of gravity. The simplest one is obtained by adding the Weyl-tensor-squared term to Starobinsky action. Such theory, called Stelle gravity [9], violates unitarity due to the presence of ghost degrees of freedom. However, they can be removed by means of a different quantization prescription that turns them into purely virtual particles [10, 11]. The modifications of Starobinsky predictions due to the presence of purely virtual particle has been derived in [12]. See also [27] for a recent review.
@@
Therefore, the condition $|\eta_{H}|\ll 1$ is violated unless $\bar{b}$ is small, which could depend on the free parameters of the theory. However, even if $\bar{b}<1$ there would be a time sufficiently far in the past where $\alpha\ll\bar{b}$, because $\alpha$ tends to zero in the infinite (conformal) past. It seems that an initial phase of the so-called constant-roll inflation [15, 28] be inevitable. We also highlight that even if we choose $\bar{b}$ small enough to mimic slow-roll inflation, the results would be quantitatively different from those of, say, class I where the linear term in the beta is absent. Indeed, the power spectra are nonanalytic in $\bar{b}=0$ and choosing a very small $\bar{b}$ could make higher-order terms larger. However, as we show below, the inverse powers of $\bar{b}$ appear at the NL and NNL order for tensor and scalar power spectra, respectively.
@@
If $\bar{b}$ is small enough we could expand (7.11) and approximate $\alpha$ as a constant. However, if we do so, the inverse powers of $\bar{b}$ that appear at higher orders could jeopardize the perturbative expansion and the reliability of the results. For these reasons this class of theories should be studied with care, in order to avoid mistakes. For example, a more appropriate treatment would be to consider the ratio $\alpha/\bar{b}$ and perform different expansions in the two limiting cases $\alpha/\bar{b}\ll 1$ and $\alpha/\bar{b}\gg 1$. We postpone a detailed study of this class to a future work. For the moment we give the power spectra at the leading order and show the NL orders assuming $\alpha$ constant to illustrate what explained above.
@@
$n_{\text{t}*}=-\frac{2\alpha_{*}^{2}\left(3-2\bar{b}\right)}{1+2\bar{b}}+\mathcal{O}(\alpha_{*}^{2}),\qquad n_{\text{s}*}-1=2\bar{b}-\frac{6\alpha_{*}\left(\bar{b}-1\right)\left(\bar{b}+b_{0}\right)}{3-2\bar{b}}+\mathcal{O}(\alpha_{*}^{2}).$ (7.22)
The tensor-to-scalar ratio at the leading order is

### v2

@@ 3.2 Scalars
We choose to work in the spatially-flat gauge, where the scalar perturbations of the metric are parametrized by
@@
Also in this case, $\sigma_{\text{s}}$ depends only on $v$ and $\beta_{\alpha}$. From the expressions of $\sigma_{\text{t,s}}$ we can derive their expansion in $\alpha$ and apply the results of [Anselmi:2021rye].
@@ 4 Predictions
The arbitrary coefficients $C_{\text{t,s}}$ can be chosen such that $w^{\text{C}}_{\text{t}}(\eta(t))=u(t)$ and $w^{\text{C}}_{\text{s}}(\eta(t))=\mathcal{R}\equiv H\frac{\delta\phi}{\dot{\bar{R}}}$, where $\mathcal{R}$ is the curvature perturbation in the spatially-flat gauge. Indeed, the quantities $w^{\text{C}}_{\text{t}}$ and $w^{\text{C}}_{\text{s}}$ can only be proportional to $u$ and $\mathcal{R}$, respectively, being gauge invariant and conserved on superhorizon scales. Comparing the redefinitions (3.24) and (3.35) with the expression (3.13) we find
@@
Then, using equations (2.17) and (2.19), the tensor-to-scalar ratio and the spectral indices read
@@
It is interesting to note that, since $r$ and $n_{\text{t,s}}$ depend only on $v$, $\beta_{\alpha}$ and $Q_{0\text{t,s}}$, they are analytic [^1] From [Anselmi:2021rye] we know that also $Q_{0\text{t,s}}$ are analytic in $b_{i}$. in $b_{i}$.
@@
These results are formally the same found in [Anselmi:2021rye] for single-field inflation in the Einstein frame.
We can now compare the results with data. The most up-to-date experimental constraints are obtained by combining the results from the Planck 2018 collaboration [Planck:2018jri], the dark energy spectroscopic instrument [DESI:2024uvr] and recent data from the Atacama Cosmology Telescope [ACT:2025fju, ACT:2025tim]. The result gives
@@
for $k_{*}=0.05\ \text{Mpc}^{-1}$. From now on every quantity with the subscript “*” is understood to be at $k=k_{*}$. From (4.13) we can derive $\alpha_{*}$ for each class and plug it back in the leading order of the tensor-to-scalar ratio $r_{*}$. In this way we obtain a relation between $r_{*}$ and $n_{\text{s}*}$. We find
@@
Using the central value for $n_{\text{s}*}$ and taking into account the bound for the tensor-to-scalar ratio $r_{*}<0.035$ from the Bicep/KECK collaboration [BICEP:2021xfz], we find that the accepted values for the first nonvanishing coefficient of the beta functions are
@@
In the literature, the number of e-folds before the end of inflation $N_{k}$ is often used as a primary constraint, since there are various theoretical estimates that place it between 50 and 60. However, those derivations depend on the inflationary model, as well as on the reheating temperature (see [Gorbunov:2011zzc] for a clear derivation of the estimate and [Zharov:2025evb] for a recent study). Moreover, the number of e-folds is also a frame-dependent quantity. For these reasons, we prefer to fix the free parameters of the models by means of the experimental data for $n_{\text{s}}$ and derive the prediction for $N_{k}$, instead of imposing $N_{k}$ in a range and then check the prediction for $n_{\text{s}}$. Then the reader can make their own interpretations of the results using the values for $N_{k}$ that they consider acceptable.
@@
If we accept that $N_{*}$ can be slightly larger than 60, then any $f(R)$ in class I is in agreement with data, provided that $b_{0}\lesssim-0.48$. On the other hand, if we impose $50\leq N_{*}\leq 60$ then class I is almost out of the two-sigma range. However, this is a mild tension and it is not enough to exclude the models in class I, considering that there are also some tensions between the data themselves [DESI:2025gwf, Zharov:2025evb]. Functions of class II are in good agreement at the one-sigma range.
@@ 5 Derivation of $f(R)$ from $\beta_{\alpha}$
In this section we follow the strategy explained in section 2 to obtain the expansions of all the background quantities for class I and II. In this way we can identify the general expression for the $f(R)$ that belong to those classes. Finally, we use the results of [Anselmi:2021rye] to work out the power spectra.
@@
Using the expression for $v$ we can write the expansions for (3.25) and (3.36), which are also analytic in $b_{i}$,
@@
From the coefficients of $\sigma_{\text{t,s}}$ we can directly obtain the solutions of the Mukhanov-Sasaki equation in the two cases using [Anselmi:2021rye]. The results up to the order $\mathcal{O}(\alpha_{k}^{3})$ for tensors and $\mathcal{O}(\alpha_{k}^{2})$ for scalars are
@@
$W_{4}$ $=$ $-\frac{16W_{0}}{1+\eta^{2}}+\frac{2(13+i\eta)W_{2}}{9(1+i\eta)}+\frac{W_{3}}{3}+4G_{2,3}^{3,1}\left(-2i\eta\left|{}_{0,0,0}^{\ 0,1}\right.\right)W_{0},$
@@ 5.1 Class I
where $F_{0}$ and $H_{0}$ are constants with $F_{0}]=0$ and $H_{0}]=1$. Using (2.21) we can write the expansion for $R$ and then invert it. We find
@@
We see that it is possible to obtain $f(R)$ with inverse and/or fractional powers of $\tilde{R}$. However, those models can only be phenomenological ones, since they do not have a well-defined Minkowski limit and/or they could not be used on every spacetime because $\tilde{R}$ could be negative in general.
In order to have more usable theories, we impose some constraints on the function. First, there should be no singularity for $\tilde{R}=0$, in order to have a well-defined Minkowski limit, as well as any other singularity for real $\tilde{R}$. Then we also avoid terms involving $\sqrt{\tilde{R}}$. These considerations force $n$ to be a positive integer [^2] A priori we could consider $n\in\mathbb{Q}$ and then fix the relations between the $b_{i}$ such that fractional powers of $R$ do not appear. However, this is not possible since the term proportional to $\tilde{R}^{-n}$ in the square bracket of (5.12) cannot be set to zero.. Such constraints are useful to obtain functions that are suitable in every spacetime background. However, if we are interested in functions that are guaranteed to be well defined around de Sitter spacetime only, we can relax those assumptions (see the end of this subsection for an example).
@@
The first is the Starobinsky model, while the second is pure quadratic gravity. In both cases a cosmological constant is present. However, in the latter it cannot be set to zero. Moreover, the $R^{2}$ term has the wrong sign and the theory propagates a ghost. This can be fixed by choosing a negative $F_{0}$, which would then lead to a negative cosmological constant. Therefore, the second model is not physically acceptable and we are left only with the Starobinsky one.
@@
Finally, if we relax one of the constraints on $f(R)$ and allow for terms containing powers of $\sqrt{\tilde{R}}$ we can have modifications to the Starobinsky model, such as
@@
where $\lambda]=0$. This model was studied in [Ketov:2010qz].
Using formulas from [Anselmi:2021rye], we obtain the following expressions for power spectra and spectral indices
@@
The coefficient $C_{\text{t,s}}$ can be obtained from (4.2) and using the explicit expressions for $F$, $H$ and $J_{\text{t,s}}$. We find
@@
This result is the same for every class. Note that the relation between $C_{\text{s}}$ and $C_{\text{t}}$ is the same as in single-field inflation with any potential [Anselmi:2021rye].
@@ 5.2 Class II
In general, the higher the class the more complicated are the singularities in the expressions of $F$ and $H$ at $\alpha=0$. Indeed, choosing $b_{0}=0$ and $b_{1}\neq 0$, the solutions of equations (2.19) are
@@
which have essential singularities. In this case inverting $R(\alpha)$ gives an expansion in powers of $1/\mathcal{W}_{0}$, where $\mathcal{W}_{0}$ is the first branch of the Lambert function, defined as
@@
[…]
@@ 6 Change of frame
In order to match the definition of $\tilde{\alpha}$ with the one used in the Einstein frame in [Anselmi:2021rye] we choose
@@
[…]
@@
which matches with the one obtained in [Anselmi:2021rye]. Moreover, the expression for $\varepsilon$ in the Einstein frame is exact and gives
@@ 7 Class 0: other polynomial $f(R)$
We conclude by discussing a special class of theories. The reader might have noticed that the simplest $f(R)$’s we can think of, i.e. polynomials, do not appear in any of the classes studied above, with the exception for the Starobinsky model and pure quadratic gravity in class I. The reason is that simple polynomial extensions of Starobinsky or Hilbert actions have a linear beta function. It is not possible to export the techniques from asymptotically-free quantum field theory and apply our RG analogy at the present stage. In particular, there are no leading logarithms that can be resummed into the running slow-roll parameter. Moreover, as discussed in section 2, in the assumption $\alpha\ll 1$ a linear beta function is sufficient to guarantee $\varepsilon\ll 1$ but leads to a constant $\eta_{H}$ [plus corrections of order $\mathcal{O}(\alpha)$] and slow-roll inflation cannot be realized. Nevertheless, this class still admits a systematic expansion in powers of $\alpha$ for the background quantities, since we only require $\alpha\ll 1$ to perturbatively solve the equations. In other words, theories with a linear beta function can still produce a phase of accelerated expansion, but the slow-roll condition $|\eta_{H}|\ll 1$ is violated, at least far enough in the past (see the discussion below).
@@
With appropriate choices of the coefficients we can obtain any polynomial extensions of the Starobinsky model, such as
@@
and so on. The same $f(R)$ without the $R^{2}$ term can also be obtained, as well as the Starobinsky model with the “wrong sign” in front of the $R^{2}$ term plus a positive cosmological constant, i.e.
@@
However, such theory propagates a scalar tachyon and is not physically acceptable.
The result just found shows an important fact about the Starobinsky model. Among all the polynomial $f(R)$ extensions of general relativity, The Starobinsky model is the only one that allows for slow-roll inflation during the initial phase of the evolution of the universe. Moreover, since all the other polynomial $f(R)$ belong to a different class, calculations cannot be obtained by suitably modifying those of the Starobinsky model, not even perturbatively. This can be seen from the fact that, like in the other classes, $F$ and $H$ are not analytic in the first nonvanishing coefficient of the beta function, in this case $\bar{b}$.
Besides their simple form, polynomials are particularly important because, in general relativity, they are generated by renormalization (together with all other possible scalar quantities built with the Riemann tensor, its covariant derivatives and the metric tensor). This means that they are always present and even if we consider them as small corrections, they cannot be viewed as such, at least in the context of slow-roll inflation. The only way to avoid their radiative generation is to consider renormalizable theories of gravity. The simplest one is obtained by adding the Weyl-tensor-squared term to Starobinsky action. Such theory, called Stelle gravity [Stelle:1976gc], violates unitarity due to the presence of ghost degrees of freedom. However, they can be removed by means of a different quantization prescription that turns them into purely virtual particles [Anselmi:2018ibi, Anselmi:2018tmf]. The modifications of Starobinsky predictions due to the presence of purely virtual particle has been derived in [Anselmi:2020lpp]. See also [Piva:2023bcf] for a recent review.
@@
Therefore, the condition $|\eta_{H}|\ll 1$ is violated unless $\bar{b}$ is small, which could depend on the free parameters of the theory. However, even if $\bar{b}<1$ there would be a time sufficiently far in the past where $\alpha\ll\bar{b}$, because $\alpha$ tends to zero in the infinite (conformal) past. It seems that an initial phase of the so-called constant-roll inflation [Motohashi:2014ppa, Motohashi:2017vdc] be inevitable. We also highlight that even if we choose $\bar{b}$ small enough to mimic slow-roll inflation, the results would be quantitatively different from those of, say, class I where the linear term in the beta is absent. Indeed, the power spectra are nonanalytic in $\bar{b}=0$ and choosing a very small $\bar{b}$ could make higher-order terms larger. However, as we show below, the inverse powers of $\bar{b}$ appear at the NL and NNL order for tensor and scalar power spectra, respectively.
@@
If $\bar{b}$ is small enough we could expand (7.11) and approximate $\alpha$ as a constant. However, if we do so, the inverse powers of $\bar{b}$ that appear at higher orders could jeopardize the perturbative expansion and the reliability of the results. For these reasons this class of theories should be studied with care, in order to avoid mistakes. For example, a more appropriate treatment would be to consider the ratio $\alpha/\bar{b}$ and perform different expansions in the two limiting cases $\alpha/\bar{b}\ll 1$ and $\alpha/\bar{b}\gg 1$. We postpone a detailed study of this class to a future work. For the moment we give the power spectra at the leading order and show the NL orders assuming $\alpha$ constant to illustrate what we explained above.
@@
$n_{\text{t}*}=-\frac{2\alpha_{*}^{2}\left(3-2\bar{b}\right)}{1+2\bar{b}}+\mathcal{O}(\alpha_{*}^{2}),\qquad n_{\text{s}*}-1=2\bar{b}-\frac{6\alpha_{*}\left(\bar{b}-1\right)\left(\bar{b}+b_{0}\right)}{3-2\bar{b}}+\mathcal{O}(\alpha_{*}^{2}),$ (7.22)
where $\psi^{(0)}(x)$ is the digamma function. The tensor-to-scalar ratio at the leading order is